# Towards Multi-agent Reinforcement Learning based Traffic Signal Control through Spatio-temporal Hypergraphs

### v1

@@ V-E Sensitive Analysis
V-E1 Impact of threshold $\zeta$
@@
V-E2 Impact of hypergraph loss coefficient $\beta$
As an essential component of our proposed algorithm, the hypergraph incorporates two types of hyperedge construction losses into the reinforcement learning loss function (see Equation 11). To investigate the impact of the hypergraph construction loss on the experimental results, we conducted a sensitivity experiment on the hypergraph loss coefficient $\beta$ in the aforementioned three datasets. The experimental results are shown in Figure 4. When the $\beta$ is set to 0.001, ATT reaches its optimal value across all three datasets. On the other hand, when $\beta$ is set to 0, it indicates that the loss incurred by hypergraph construction is ignored, resulting in substantial decrease in ATT performance compared to the optimal value.
V-E3 Impact of Attention Head Number $K$
@@ VI Conclusion
In this paper, we propose a framework that can cooperate with multiple edge computing servers for realizing intelligent traffic signal control. Within this framework, we introduce a multi-agent collaborative reinforcement learning algorithm based on MA-SAC in multi-intersection environments. In order to promote the effective decision-making and improved processing of spatio-temporal information among multiple agents, we incorporate hypergraph learning into the MA-SAC. We design the dynamic generation of spatio-temporal hyperedges to construct hypergraphs with enhanced modeling capabilities. We then utilize multi-head attention to update intersection information, enabling dynamic spatio-temporal interactions between road intersections. The results obtained from extensive experiments demonstrate that our proposed method outperforms state-of-the-art ones in traffic signal control, achieving improvements in both travel time and throughput. Our future direction is to study data cleaning and integration schemes to mitigate the impact of missing or noisy existing in the collected input data, which aims to enhance the robustness of our proposed method.

### v2

@@ V-F Sensitive Analysis
V-F1 Impact of threshold $\zeta$
@@
V-F2 Impact of hypergraph loss coefficient $\beta$
As an essential component of our proposed method, the hypergraph incorporates two types of hyperedge construction losses into the reinforcement learning loss function (see Equation 12). To investigate the impact of the hypergraph construction loss on the experimental results, we conducted a sensitivity experiment on the hypergraph loss coefficient $\beta$ in the aforementioned three datasets. The selected range [0, 0.1] for the $\beta$ value is chosen to effectively scale the two components of the loss function (see Equation 12). This ensures their values fall within a comparable range, facilitating a thorough analysis of the trade-off between them. The experimental results are shown in Figure 6. When $\beta$ is set to 0.001, ATT reaches its optimal value across all three datasets. This observation confirms that satisfactory results are not solely dependent on a single $\beta$ value in our proposed method, but rather occur within a consistent range of $\beta$ values. On the other hand, when $\beta$ is set to 0, it indicates that the loss incurred by hypergraph construction is ignored, resulting in substantial decrease in ATT performance compared to the optimal value. While the results demonstrate the promising performance of our proposed method in the current experimental setup, achieving robust sensitivity performance remains a challenging task, which should be further investigated in future work.
V-F3 Impact of Attention Head Number $K$
@@ VI Conclusion
In this paper, we propose a framework that can cooperate with multiple edge computing servers for realizing intelligent traffic signal control. Within this framework, we introduce a multi-agent collaborative reinforcement learning algorithm based on MA-SAC in multi-intersection environments. In order to promote the effective decision-making and improved processing of spatio-temporal information among multiple agents, we incorporate hypergraph learning into the MA-SAC. We design the dynamic generation of spatio-temporal hyperedges to construct hypergraphs with enhanced modeling capabilities. We then utilize multi-head attention to update intersection information, enabling dynamic spatio-temporal interactions between road intersections. The results obtained from extensive experiments demonstrate that our proposed method outperforms state-of-the-art ones in traffic signal control, achieving improvements in both travel time and throughput, while exhibiting a more stable post-convergence performance. Our future direction is to study data cleaning and integration schemes to mitigate the impact of missing or noisy existing in the collected input data, which aims to enhance the robustness of our proposed method.